OpenStudy (anonymous):

natural log of 10000*exp(x)

OpenStudy (anonymous):

when you take the natural log of \[10000*e^x\]equal to \[\ln(10000)*x\]

OpenStudy (anonymous):

@zepdrix can you double check if my memory is correct? Having trouble confirming this law on Google

OpenStudy (anonymous):

is it the product rule?\[\ln(xy)=\ln x + \ln y\]

OpenStudy (anonymous):

so it would be\[\ln 10000 + x\]

zepdrix (zepdrix):

Your product rule looks correct! Ya ya there ya go, addition between them!

OpenStudy (anonymous):

Thank you thank you